Solve the differential equation:
(D^4-a^4)y=x^4+sinbx, where a and b are constants and D=d/dx
1
Expert's answer
2020-11-30T21:07:01-0500
"Solution"

It's AE is "D^4-a^4=0"

"\therefore\ the\ auxiliary\ equation\ becomes\ (D^2 – a^2 ) (D^2 + a^2 ) = 0\\
\implies D^2-a^2=0\ or\ D^2+a^2=0\\
\therefore (D – a) (D + a) (D – ai) (D + ai) = 0\\
\therefore D = \pm a, and\ D=\pm ai"


Hence the required solution is

"C.F=[ae^{ax}+be^{-ax}]+[ce^{iax}+de^{-iax}]\\
C.F =[a \frac{(cosh\ ax+sinh\ ax)}{2}]+[b \frac{(cosh\ ax-sinh\ ax)}{2}]+[c \frac{(cos\ ax-i\ sin\ ax)}{2}+d \frac{(cos\ ax-sin\ ax)}{2}]\\
C.F=[\frac{a+b}{2}cos\ ax + \frac{a-b}{2}sinh\ ax]+[\frac{c+d}{2}cos\ ax + i\frac{(c-d)}{2}sin\ ax]"


Where "C_1=\frac{a+b}{2},\ C_2=\frac{a-b}{2},\ C_3=\frac{c+d}{2},\ c_4=\frac{i(c-d)}{2},\"

Its solution is


"C.F=[C_1\ cosh\ ax +C_2\ sinh\ ax]+[C_3\ cos\ ax +C_4\ sin\ ax]---(i)"

And PI, for "x^4" is;


"PI_1=\frac{1}{D^4-a^4}\ x^4"

Putting "D^4=a^4" in "f(D)", then

"PI_1=\frac10 x^4\ \therefore method\ fails"


Then to find PI, put "D^4+a^4" for "f(D)" and take out "x^4".

"\therefore PI_1=x^4 \{\frac{1}{D^4-a^4+a^4}\}\{1\}\\
\implies x^4\ \frac{1}{D^4}\{1\}=x^4\frac{x^4}{4!} \implies \frac{x^8}{4!}-----(ii)"


Also, "PI_2" for "sin\ bx" is;


"\therefore PI_2=\frac{1}{D^4-a^4}\ sin\ bx\\"

Put "D^4=D^2 \cdot D^2 = (-b^2)(-b^2)"

"\implies\frac{1}{(-b^2)(-b^2)-a^4}\ sin\ bx = \frac{1}{b^4-a^4}\ sin\ bx ---(iii)"


General Solution is;

"G.S=C.F+PI_1+PI_2 \implies [C_1\ cosh\ ax +C_2\ sinh\ ax]+[C_3\ cos\ ax +C_4\ sin\ ax]+\frac{x^8}{4!}+\frac{1}{b^4-a^4}\ sin\ bx"


Hence

"G.S=[C_1\ cosh\ ax +C_2\ sinh\ ax]+[C_3\ cos\ ax +C_4\ sin\ ax]+\frac{x^8}{4!}+\frac{1}{b^4-a^4}\ sin\ bx"